OpenStudy (anonymous):

If -nx2 + tx + c = 0, what is x equal to? x is equal to minus t plus or minus square root of t squared minus 4 multiplied by n multiplied by c, all divided by minus 2 times n x is equal to t plus or minus square root of t squared plus 4 multiplied by n multiplied by c, all divided by minus 2 times n x is equal to t plus or minus square root of t squared minus 4 multiplied by n multiplied by c, all divided by minus 2 times n x is equal to minus t plus or minus square root of t squared plus 4 multiplied by n multiplied by c, all divided by minus 2 times n

OpenStudy (anonymous):

The answer is A

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

x=-t+-square root t^2-4ac/-2n is that correct?

OpenStudy (phi):

it's hard to read equations written out in words. but the formula to find x so that \[ a x^2 + bx + c=0 \] is \[ \frac{1}{2a} \left(-b \pm \sqrt{b^2 -4ac} \right)\]

OpenStudy (phi):

be careful about the signs. what is a, b and c in your problem ?

OpenStudy (anonymous):

that the thing idk

OpenStudy (anonymous):

i just looked at the quadratic formula and that is it.

OpenStudy (phi):

\[ -nx^2 + tx + c = 0 \] what is "a" ?

OpenStudy (anonymous):

-n

OpenStudy (phi):

ok, and b is t and c is c use a= -n, b= t and c=c in the quadratic formula.

OpenStudy (anonymous):

ohh so its D

OpenStudy (anonymous):

x=-t+-square root t^2+4ac/-2n

OpenStudy (anonymous):

4nc i ment

OpenStudy (phi):

yes

OpenStudy (phi):

the - 4(-n)c becomes +4nc

OpenStudy (phi):

if you see -x^2 -2x -3 that means -1*x^2 + -2*x + -3 and a = -1, b=-2 and c= -3